Find the equation of the cone with the vertex at (1,−1,2) and the base curve as
(z+1)^2 = x+2, y = 3.
1
Expert's answer
2020-05-15T15:54:20-0400

Solution:

The vertex "V(1,-1,2)". Let "a,b,c" be the direction ratios of generator of the cone.

Then the equations of the generator are

"\frac{x-1}{a}=\frac{y+1}{b}=\frac{z-2}{c}=t" (say) (I)

The coordinates of any points on the generator are "(1+at,-1+bt,2+ct)".

For some "t\in \R, (1+at,-1+bt,2+ct)" lies on the guiding curve.

Therefore "((2+ct)+1)^2=(1+at)+2" and "-1+bt=3" . Thus, "t=\frac{4}{b}".

From this we get

"((2+c*\frac{4}{b})+1)^2=(1+a*\frac{4}{b})+2"

From (I) we obtain:

"(3+4*\frac{z-2}{y+1})^2=3+4*\frac{x-1}{y+1}"

After simplification, the required equation of the cone is

"6y^2+24zy-4xy-32y+16z^2+26-40z-4x=0."

Answer:

"6y^2+24zy-4xy-32y+16z^2+26-40z-4x=0."